OpenStudy (anonymous):

How do you write an equation of a line so that it is parallel or perpendicular to a given point and a given line?

OpenStudy (anonymous):

can yo help me please i really need to get this done

OpenStudy (tukitw):

You probably meant at a given point because it's impossible to have a line parallel or perpendicular to a point

OpenStudy (tukitw):

Suppose the given point is (x1 , y1). The typical formula for finding an equation of the line would be y - y1 = m(x - x1). For this line to be parallel with the given line, m will be the same as the gradient of that given line, which you can read easily from the form of y = nx + c, where n is the gradient. So, just substitute m = n and the given coordinates to get the equation of the parallel line. Likewise for a perpendicular line, m will be the same as the negative reciprocal of the gradient of that given line. So, substitute m = -1 / n and the given coordinates to get the equation of the perpendicular line.